OpenStudy (anonymous):

a ball bounces from a height of 2 metres and returns to 80% of its previous height on each bounce. find the total distance travelled by the ball until it stops bouncing a ball bounces from a height of 2 metres and returns to 80% of its previous height on each bounce. find the total distance travelled by the ball until it stops bouncing @Mathematics

OpenStudy (anonymous):

use the sum for geometric series equation

OpenStudy (anonymous):

ok.. i know but how do you do that?

OpenStudy (anonymous):

sum=18meters

OpenStudy (anonymous):

2/(1-0.8) = 2/0.2 = 2*5 = 10m

OpenStudy (anonymous):

you use the formula sum to infinity = a / (1 - r) where a = first term (2) and r = commom ratio (0.8) - as jamesm did.

OpenStudy (phi):

Here's more detail: So for the moment ignore the initial drop of 2 m. After that we have pairs up and down. Find just the down (and then double to account for the up): with d= 2m (the original height) we have 0.8d+ 0.8 0.8 d + 0.8 0.8 0.8 d + ... or d 0.8 (1+0.8+0.8^2 + ...) the series (with a= 1, r= 0.8) sums to a/(1-r)= 1/(1-.8)= 1/.2= 5 d 0.8 5= 4d= 8m That's the down. Double it for up and down, 16 m add the first drop of 2m and finally we get 18m